OpenStudy (anonymous):

does y' of ln(4^s) = 1/4^s or s/4 ??

OpenStudy (anonymous):

\(\large ln(4^s) = s \cdot ln(4)\)

OpenStudy (anonymous):

so it would be s/4 then.

OpenStudy (anonymous):

ln(4) is a constant

OpenStudy (anonymous):

yes but the derivative of ln=1/x

OpenStudy (anonymous):

If you are finding \(\large y'=\frac{d}{ds}[ ln(4^s)]\) \(\large y'=\frac{d}{ds}[ ln(4)s] =\frac{d}{ds}[c\cdot s], \space c=ln(4)\)

OpenStudy (anonymous):

The derivative of ln x is 1/x, but ln 4 is a constant, and the derivative of a constant is zero.

OpenStudy (anonymous):

But here you are taking the derivative of a constant times a variable.

OpenStudy (anonymous):

what is c*s ???

OpenStudy (anonymous):

im confused by your explanation.

OpenStudy (turingtest):

the derivative is with respect to s what is d/ds(s) ?

OpenStudy (anonymous):

umm s'?

OpenStudy (turingtest):

no, what is d/dx(x) ?

OpenStudy (anonymous):

s' would be if s was a function of some other variable.

OpenStudy (anonymous):

=1

OpenStudy (turingtest):

yes, so we just changed the name of s to x ...so d/ds(s)=?

OpenStudy (anonymous):

=1 !

OpenStudy (turingtest):

yes :) now what if we multiply by a constant that we call 'c' ? then d/ds(cs)=?

OpenStudy (anonymous):

c

OpenStudy (turingtest):

yes now for your problem we can write ln(4^s)=s*ln4 ln4 is a constant, so d/ds(ln(4^s))=?

OpenStudy (anonymous):

ln4?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

so thats the answer?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

Thanks for the backup, Turing!

OpenStudy (turingtest):

No prob!